# Dealing with unbounded gradients in stochastic saddle-point optimization

### v1

@@ 1 Introduction
and primal-dual gradient descent starting from the initial point $\bm{x}_{1}=0$ and $\bm{y}=0$ as a baseline:
$\bm{x}_{t+1}$ $=\bm{x}_{t}-\eta\nabla_{x}f(\bm{x}_{t},\bm{y}_{t})=\bm{x}_{t}-\eta\bm{g}_{x}(t)$
$\bm{y}_{t+1}$ $=\bm{y}_{t}-\eta\nabla_{y}f(\bm{x}_{t},\bm{y}_{t})=\bm{y}_{t}-\eta\bm{g}_{y}(t),$
where we also introduced the shorthand notations $\bm{g}_{x}(t)=\nabla_{x}f(\bm{x}_{t},\bm{y}_{t})$ and $\bm{g}_{y}(t)=\nabla_{y}f(\bm{x}_{t},\bm{y}_{t})$. Using standard tools (that we will explain in detail below), the average of the first $T$ iterates produced by the above procedure can be shown to satisfy the following guarantee on the duality gap:
@@
$\quad\leq\frac{\left\|\bm{x}^{*}\right\|_{2}^{2}+\left\|\bm{y}^{*}\right\|_{2}^{2}}{2\eta T}+\frac{\eta}{2T}\sum_{t=1}^{T}\left(\left\|\bm{g}_{x}(t)\right\|_{2}^{2}+\left\|\bm{g}_{y}(t)\right\|_{2}^{2}\right).$
If one can ensure that the gradients $\bm{g}_{x}(t)$ and $\bm{g}_{y}(t)$ remain bounded by a constant $G>0$, one can set $\eta\sim 1/(G\sqrt{T})$ and obtain a convergence rate of order $\frac{G\left(\left\|\bm{x}^{*}\right\|_{2}^{2}+\left\|\bm{y}^{*}\right\|_{2}^{2}\right)}{\sqrt{T}}$. However, notice that there is no way to make sure that the gradients actually remain bounded while executing the algorithm! Indeed, notice that $\nabla_{x}f(\bm{x}_{t},\bm{y}_{t})=\bm{M}\bm{y}_{t}+\bm{b}$, which grows large as $\bm{y}_{t}$ grows large. A natural idea is to project the iterates to balls of respective sizes $D_{x},\,D_{y}>0$, which guarantees that the iterates and thus the gradients remain bounded. However, convergence to the saddle point $(\bm{x}^{*},\bm{y}^{*})$ is now only possible whenever the respective norms satisfy $\left\|\bm{x}^{*}\right\|\leq D_{x}$ and $\left\|\bm{y}^{*}\right\|\leq D_{y}$, otherwise the optimal solution is excluded from the feasible set. Unfortunately, in many applications, it is impossible to pick the constants $D_{x}$ and $D_{y}$ appropriately due of lack of prior knowledge of the solution norms. We give an important example of such a situation at the end of this section.
@@
where the big-O notation hides some problem-dependent constants related to the objective function $f$ (which will be made explicit in our main theorem). Notably, our method requires no prior knowledge of the norms $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}$ and $\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}$ whatsoever, and in particular performs no projections to make sure that the iterates remain bounded, thus addressing the challenge outlined above. As we will show, this guarantee holds even when the gradients are subject to multiplicative noise that can scale with the magnitude of the gradient itself. Our main technical tool is augmenting the objective with a well-chosen regularization term which allows us to eliminate the terms $\left\|\bm{g}_{x}(t)\right\|_{2}^{2}+\left\|\bm{g}_{y}(t)\right\|_{2}^{2}$ appearing in the guarantee of standard primal-dual gradient descent, and replace them with an upper bound of the gradients of the objective evaluated at the initial point $(\bm{x}_{1},\bm{y}_{1})$. These bounds have the appealing property of being initialization-dependent, in that they guarantee improved performance when we pick the initial points $(\bm{x}_{1},\bm{y}_{1})$ close to $(\bm{x}^{*},\bm{y}^{*})$.
Our contributions are closely related to the work of Liu & Orabona (2022), who propose algorithms that are guaranteed to achieve an initialization-dependent convergence rate of the order $G(\bm{x}^{*},\bm{y}^{*})=\widetilde{\mathcal{O}}\left(\frac{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}}{\sqrt{T}}\right)$, where $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors of $T$ and the solution norms. When compared to our main result, this bound demonstrates an improved scaling with the initialization error. However, their result is proved under the condition that all gradients remain bounded: $\left\|\bm{g}_{x}(t)\right\|_{2}^{2}\leq 1$ and $\left\|\bm{g}_{y}(t)\right\|_{2}^{2}\leq 1$. As per the above discussion, this is only possible in general when projecting the iterates to a bounded subset of the domain, which requires prior knowledge of the norms $\left\|\bm{x}^{*}\right\|$ and $\left\|\bm{y}^{*}\right\|$. When accounting for this issue and adapting the method of Liu & Orabona (2022) to our setting, their bounds end up scaling with $\left(D_{x}+D_{y}\right)\left(\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}\right)$, which is worse than the scaling of our bound. Furthermore, the results of Liu & Orabona (2022) are only proved for noiseless gradients (although we believe that this specific restriction may not be hard to address). On the positive side, this assumption allows their algorithm to achieve very fast convergence when initialized very close to the optimum—which is generally not possible in the stochastic case we consider. Taking all this into account, we regard our guarantees as significantly different from their results, and in many senses a strict improvement.
@@
More broadly speaking, our work contributes to the line of work on parameter-free optimization methods that are able to adapt to problem complexity without prior knowledge of the relevant problem parameters. In the context of online convex optimization (OCO), several effective parameter-free algorithms are known to achieve guarantees scaling optimally with the initialization error $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|$, without requiring prior knowledge thereof (Streeter & McMahan, 2012; Orabona, 2013; van der Hoeven, 2019). (Cutkosky & Boahen, 2016; Cutkosky, 2019; Mhammedi & Koolen, 2020) improve these guarantees by providing initialization-adaptive bounds for OCO in unconstrained domains without prior knowledge of both the size of the domain or sub-gradients of the loss. One would think that this would make their method suitable for solving the problem we study in this paper—however, their bound depends on the maximum norm of the observed sub-gradients, which is problematic for the reasons we have discussed extensively above.
@@
Unconstrained saddle-point problems have many important applications. Perhaps the most well-known such applications is in optimizing dual representations of convex functions (Bubeck et al., 2015, Section 5.2,Shalev-Shwartz & Singer, 2006; Wang & Abernethy, 2018; Wang et al., 2023). Our original motivation during the development of this work has been to develop primal-dual methods for solving average-reward Markov decision processes (MDPs): this problem can be formulated as a linear program with primal variables that are of unknown scale. In the simpler setting of discounted Markov decision processes, previous work has provided efficient planning methods based on saddle-point optimization (Wang, 2017; Jin & Sidford, 2020; Cheng et al., 2020). While in this simple setting the primal variables (called value functions in this setting) are known to be uniformly bounded, this is not the case in the more challenging average-reward setting we consider here: in this case, the value functions can have arbitrarily large norm depending on the program structure. As it is well-known in the reinforcement-learning literature, estimating this parameter is as hard as solving the original problem, and learning optimal policies without its prior knowledge has been widely conjectured to be impossible (Bartlett & Tewari, 2009; Fruit et al., 2018b, a; Zhang & Ji, 2019). Using our techniques developed in the present paper, we make progress on this important problem by proposing a planning algorithm that is guaranteed to produce a near-optimal policy without having prior knowledge of the scale of the value functions after a polynomial number of queries made to a simulator of the environment.
@@ 2 Preliminaries
The final output of the algorithm will be denoted as $(\overline{\bm{x}}_{T},\overline{\bm{y}}_{T})$, and due to noise in the gradients, its quality will be measured in terms of the expected duality gap, defined with respect to a comparator point $(\bm{x}^{*},\bm{y}^{*})$ as
$\mathbb{E}\left[G(\bm{x}^{*},\bm{y}^{*})\right]=\mathbb{E}\left[f(\overline{\bm{x}}_{T},\bm{y}^{*})-f(\bm{x}^{*},\overline{\bm{y}}_{T})\right].$
Here, it is typical to choose as comparator point a saddle point of $f$ that satisfies the inequalities
$f(\bm{x}^{*},\bm{y})\leq f(\bm{x}^{*},\bm{y}^{*})\leq f(\bm{x},\bm{y}^{*}).$ (2)
for all $\bm{x}\in\mathcal{X},\bm{y}\in\mathcal{Y}$. Our analysis will allow proving guarantees against arbitrary comparator points, which is useful for certain applications (cf. Section 4).
@@ 3.1 Unconstrained bilinear games
Let $\varrho_{y}=2\eta_{x}L_{M}^{2}$ and $\varrho_{x}=2\eta_{y}L_{M}^{2}$. Then, the duality gap achieved by COGDA satisfies the following bound against any comparator $(\bm{x}^{*},\bm{y}^{*})\in^{m}\times^{n}$:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{2\eta_{y}T}+\eta_{x}L_{M}^{2}\right)\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}$
$\quad+\left(\frac{1}{2\eta_{x}T}+\eta_{y}L_{M}^{2}\right)\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}$
$\quad+\frac{\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\frac{\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right]$
In particular, setting $\bm{x}_{1}=0$ and $\bm{y}_{1}=0$ and $\eta_{x}=1/{L_{M}\sqrt{T}}$ and $\eta_{y}=1/{L_{M}\sqrt{T}}$, the duality gap is upper bounded as
$\mathbb{E}\left[G(\bm{x}^{*},\bm{y}^{*})\right]$ $\!=\!\mathcal{O}\left(\frac{L_{M}^{2}\left(\left\|\bm{y}^{*}\right\|_{2}^{2}+\left\|\bm{x}^{*}\right\|_{2}^{2}\right)+L_{b}^{2}+L_{c}^{2}}{L_{M}\sqrt{T}}\right).$
It is insightful to compare this bound side by side with the one we would get by running primal-dual stochastic gradient without regularization. By standard arguments (see, e.g., Liu & Orabona, 2022; Abernethy et al., 2018; Zinkevich, 2003), the following bound is easy to prove:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]\leq$ $\frac{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}}{\eta_{x}T}+\frac{\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}}{2\eta_{y}T}$
$+\frac{\eta_{x}}{2T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{t}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right]$
$+\frac{\eta_{y}}{2T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{t}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right].$
@@
At first, it may seem surprising that such an improvement is possible to achieve by such a simple regularization trick. To provide some insight about how regularization helps us achieve our goal, we provide the brief proof sketch of the above statement here.
@@
Fix $(\bm{x}^{*},\bm{y}^{*})\in^{m}\times^{n}$. As the first step, we introduce the notation $f^{\left(\text{reg}\right)}\left(\bm{x},\bm{y}\right)=f\left(\bm{x},\bm{y}\right)+\frac{\varrho_{x}}{2}\left\|\bm{x}-\bm{x}_{1}\right\|_{2}^{2}-\frac{\varrho_{y}}{2}\left\|\bm{y}-\bm{y}_{1}\right\|_{2}^{2}$ and rewrite the expected duality gap as
@@
These terms can then be controlled via the standard regret analysis of COMID due to Duchi et al. (2010). In particular, a few lines of calculations (along the lines of the online gradient descent analysis of Zinkevich, 2003) yield the following bound on the sum of the two regrets:
@@
$\qquad\leq\frac{\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}}{2\eta_{y}T}+\frac{\eta_{y}}{2T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{2}^{2}\right]$
$\qquad\quad+\frac{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}}{2\eta_{x}T}+\frac{\eta_{x}}{2T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{2}^{2}\right]$
@@
As can be seen from the proof, the role of the additional regularization term for the $x$-player is to eliminate the gradient norms appearing in the regret bound of the $y$-player. This effect kicks in once the regularization parameter $\varrho_{x}$ becomes large enough, so that the corresponding negative term in the regret bound of the first player can overpower the positive term appearing on the bound of the opposite player. The same story applies to the second player. Note that while the regularization pulls the iterates closer to the initial point $\bm{x}_{1},\bm{y}_{1}$, it does not explicitly guarantee that they remain uniformly bounded at all times $t$, and in fact such claim seems impossible to show in general due to the noise in the gradient estimates. Remarkably, the analysis above works seamlessly for noisy gradient estimates, even though the gradient noise can grow proportionally with the size of the iterates.

### v2

@@ 1 Introduction
and primal-dual stochastic gradient descent ascent starting from the initial point $\bm{x}_{1}=0$ and $\bm{y}=0$ as a baseline:
$\bm{x}_{t+1}$ $=\bm{x}_{t}-\eta\widetilde{\bm{g}}_{x}(t)$
$\bm{y}_{t+1}$ $=\bm{y}_{t}-\eta\widetilde{\bm{g}}_{y}(t),$
where $\widetilde{\bm{g}}_{x}(t)=\nabla_{x}f(\bm{x}_{t},\bm{y}_{t})+\xi_{x}(t)$ and $\widetilde{\bm{g}}_{y}(t)=\nabla_{y}f(\bm{x}_{t},\bm{y}_{t})+\xi_{y}(t)$ are potentially noisy and unbiased estimates of the gradients with respect to $\bm{x}$ and $\bm{y}$. Using standard tools (that we will explain in detail below), the average of the first $T$ iterates produced by the above procedure can be shown to satisfy the following guarantee on the duality gap:
@@
$\,\leq\frac{\left\|\bm{x}^{*}\right\|_{2}^{2}+\left\|\bm{y}^{*}\right\|_{2}^{2}}{2\eta T}+\frac{\eta}{2T}\sum_{t=1}^{T}{\mathbb{E}\left[\left\|\widetilde{\bm{g}}_{x}(t)\right\|_{2}^{2}+\left\|\widetilde{\bm{g}}_{y}(t)\right\|_{2}^{2}\right]}.$
If one can ensure that the gradient estimates $\widetilde{\bm{g}}_{x}(t)$ and $\widetilde{\bm{g}}_{y}(t)$ remain bounded by a constant $G>0$, one can set $\eta\sim 1/(G\sqrt{T})$ and obtain a convergence rate of order $\frac{G\left(\left\|\bm{x}^{*}\right\|_{2}^{2}+\left\|\bm{y}^{*}\right\|_{2}^{2}\right)}{\sqrt{T}}$. However, notice that there is no way to make sure that the gradients actually remain bounded while executing the algorithm! Indeed, notice that $\nabla_{x}f(\bm{x}_{t},\bm{y}_{t})=\bm{M}\bm{y}_{t}+\bm{b}$, which grows large as $\bm{y}_{t}$ grows large. A natural idea is to project the iterates to balls of respective sizes $D_{x},\,D_{y}>0$, which guarantees that the iterates and thus the gradients remain bounded. However, convergence to the saddle point $(\bm{x}^{*},\bm{y}^{*})$ is now only possible whenever the respective norms satisfy $\left\|\bm{x}^{*}\right\|\leq D_{x}$ and $\left\|\bm{y}^{*}\right\|\leq D_{y}$, otherwise the optimal solution is excluded from the feasible set. Unfortunately, in many applications, it is impossible to pick the constants $D_{x}$ and $D_{y}$ appropriately due of lack of prior knowledge of the solution norms.
@@
where the big-O notation hides some problem-dependent constants related to the objective function $f$ (which will be made explicit in our main theorem). Notably, our method requires no prior knowledge of the norms $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}$ and $\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}$ whatsoever, and in particular performs no projections to make sure that the iterates remain bounded, thus addressing the challenge outlined above. Furthermore, our guarantees continue to hold for potentially data-dependent comparators $(\bm{x}^{*},\bm{y}^{*})$, and even when the gradients are subject to multiplicative noise that can scale with the magnitude of the gradient itself. Our main technical tool is augmenting the objective with a well-chosen regularization term which allows us to eliminate the terms $\left\|\widetilde{\bm{g}}_{x}(t)\right\|_{2}^{2}+\left\|\widetilde{\bm{g}}_{y}(t)\right\|_{2}^{2}$ appearing in the guarantee of standard primal-dual gradient descent, and replace them with an upper bound of the gradients of the objective evaluated at the initial point $(\bm{x}_{1},\bm{y}_{1})$. These bounds have the appealing property of being initialization-dependent, in that they guarantee improved performance when we pick the initial points $(\bm{x}_{1},\bm{y}_{1})$ close to $(\bm{x}^{*},\bm{y}^{*})$. Besides the simple bilinear setting discussed above, we study a more general set of games that we call “sub-bilinear”, and provide an algorithmic framework that comes with guarantees that express closeness to initialization in terms of general Bregman divergences.
The initialization-adaptive nature of our bounds is similar to the guarantees proved by Liu & Orabona (2022), who propose algorithms for the same setting that achieve an initialization-dependent convergence rate of the order $G(\bm{x}^{*},\bm{y}^{*})=\widetilde{\mathcal{O}}\left(\frac{\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}+\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}}{\sqrt{T}}\right)$, where $\widetilde{\mathcal{O}}(\cdot)$ hides polylogarithmic factors of $T$ and the comparator norms. While this guarantee may appear to be stronger than our result, it is only proved under the condition that all gradients remain bounded as $\left\|\bm{g}_{x}(t)\right\|_{2}^{2}\leq 1$ and $\left\|\bm{g}_{y}(t)\right\|_{2}^{2}\leq 1$, which cannot be ensured in the unconstrained setting that we consider in the present work. Another closely related work (that was pointed to our attention after posting the initial version of this work online) is due to Jacobsen & Cutkosky (2023), who proposed an algorithm for convex optimization that is able to deal with potentially unbounded gradients that satisfy a certain “sub-quadratic” growth condition. As their Section 3 shows, their technique is directly extensible to saddle-point optimization in bilinear games and even our notion of sub-bilinear games. Even more curiously, their key technical idea is nearly identical to the one employed in our work. However, the objectives of their work were slightly different from ours in that they focused on the special case of Euclidean geometries and deterministic gradients. In contrast, our main results hold for stochastic gradients and stochastic comparator points $(\bm{x}^{*},\bm{y}^{*})$, and are expressed in terms of general Bregman divergences. We believe that their analysis can be adjusted to handle these extensions: while dealing with stochastic gradients seems easy, handling data-dependent comparator points and working with general Bregman divergences appears to be rather challenging in their framework, and could require significant changes to their algorithm and analysis. We nevertheless encourage the reader to credit Jacobsen & Cutkosky (2023) just as much as our work when referring to the regularization technique we study in this paper.
@@
More broadly speaking, our work contributes to the line of work on parameter-free optimization methods that are able to adapt to problem complexity without prior knowledge of the relevant problem parameters. In the context of online convex optimization (OCO), several effective parameter-free algorithms are known to achieve guarantees scaling optimally with the initialization error $\left\|\bm{x}^{*}-\bm{x}_{1}\right\|$, without requiring prior knowledge thereof (Streeter & McMahan, 2012; Orabona, 2013; van der Hoeven, 2019). Cutkosky & Boahen (2016); Cutkosky (2019); Mhammedi & Koolen (2020) improve these guarantees by providing initialization-adaptive bounds for OCO in unconstrained domains without prior knowledge of both the size of the domain or subgradients of the loss. One would think that this would make their method suitable for solving the problem we study in this paper—however, their bound depends on the maximum norm of the observed subgradients, which is problematic for the reasons we have discussed extensively above.
@@
Unconstrained saddle-point problems have many important applications. Perhaps the most well-known such applications are in optimizing dual representations of convex functions (Shalev-Shwartz & Singer, 2006; Bubeck et al., 2015, Section 5.2;Wang & Abernethy, 2018; Wang et al., 2023). Our original motivation during the development of this work has been to develop primal-dual methods for solving average-reward Markov decision processes (MDPs): this problem can be formulated as a linear program with primal variables that are of unknown scale. In the simpler setting of discounted Markov decision processes, previous work has provided efficient planning methods based on saddle-point optimization (Wang, 2017; Jin & Sidford, 2020; Cheng et al., 2020). While in this simple setting the primal variables (called value functions in this setting) are known to be uniformly bounded, this is not the case in the more challenging average-reward setting we consider here: in this case, the value functions can have arbitrarily large norm depending on the program structure. As it is well-known in the reinforcement-learning literature, estimating this parameter with only online sample access is as hard as solving the original problem, and learning optimal policies without its prior knowledge has been widely conjectured to be impossible (Bartlett & Tewari, 2009; Fruit et al., 2018b, a; Zhang & Ji, 2019). Using our techniques developed in the present paper, we make progress on this important problem by proposing a planning algorithm that is guaranteed to produce a near-optimal policy without having prior knowledge of the scale of the value functions after a polynomial number of queries made to a simulator of the environment.
@@ 2 Preliminaries
The final output of the algorithm will be denoted as $(\overline{\bm{x}}_{T},\overline{\bm{y}}_{T})$, and due to noise in the gradients, its quality will be measured in terms of the duality gap, defined with respect to a comparator point $(\bm{x}^{*},\bm{y}^{*})$ as
$G(\bm{x}^{*},\bm{y}^{*})=f(\overline{\bm{x}}_{T},\bm{y}^{*})-f(\bm{x}^{*},\overline{\bm{y}}_{T}).$
We will allow the comparator points $\bm{x}^{*},\bm{y}^{*}$ to depend on the entire interaction history $\mathcal{F}_{T}$, and in particular allow choices such as $(\bm{x}^{*},\bm{y}^{*})=\mathop{\mbox{ arg\,max}}_{\bm{x},\bm{y}\in\mathcal{S}}G(\bm{x},\bm{y})$ for arbitrary bounded sets $\mathcal{S}\subseteq\mathcal{X}\times\mathcal{Y}$. The gap function evaluated at this point is often called a merit function, which measures progress towards finding an optimal solution to the min-max optimization problem in question. Another typical choice for comparator point is a saddle point of $f$ that satisfies the inequalities
$f(\bm{x}^{*},\bm{y})\leq f(\bm{x}^{*},\bm{y}^{*})\leq f(\bm{x},\bm{y}^{*})$ (2)
for all $\bm{x}\in\mathcal{X},\bm{y}\in\mathcal{Y}$. In the rest of the paper, we will use $\bm{x}^{*},\bm{y}^{*}$ to refer more generally to any potentially random pair of comparator points, and will call such comparator choices adaptive. We provide an example that makes use of adaptive comparators in Section 4 in the context of reinforcement learning.
@@ 3.1 Unconstrained bilinear games
Let $\varrho_{y}=4\eta_{x}L_{M}^{2}$ and $\varrho_{x}=4\eta_{y}L_{M}^{2}$. Then, the duality gap achieved by COGDA satisfies the following bound against any adaptive comparator $(\bm{x}^{*},\bm{y}^{*})\in^{m}\times^{n}$:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]$ $\leq\left(\frac{1}{\eta_{y}T}+2\eta_{x}L_{M}^{2}\right)\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]$
$\quad+\left(\frac{1}{\eta_{x}T}+2\eta_{y}L_{M}^{2}\right)\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]$
$\quad+\frac{2\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{1}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right]$
$\quad+\frac{2\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{1}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right].$
In particular, setting $\bm{x}_{1}=0$ and $\bm{y}_{1}=0$ and $\eta_{x}=1/{L_{M}\sqrt{2T}}$ and $\eta_{y}=1/{L_{M}\sqrt{2T}}$, the duality gap is upper bounded as
$\mathbb{E}\left[G(\bm{x}^{*},\bm{y}^{*})\right]$ $\!=\!\mathcal{O}\left(\frac{L_{M}^{2}\mathbb{E}\left[\left\|\bm{y}^{*}\right\|_{2}^{2}+\left\|\bm{x}^{*}\right\|_{2}^{2}\right]+L_{b}^{2}+L_{c}^{2}}{L_{M}\sqrt{T}}\right).$
We stress that this bound holds against arbitrary data-dependent choices of the comparator points $(\bm{x}^{*},\bm{y}^{*})$, and in particular also against the choice $(\bm{x}^{*},\bm{y}^{*})=\mathop{\mbox{ arg\,max}}_{\bm{x},\bm{y}\in\mathcal{S}}G(\bm{x},\bm{y})$ for any bounded set $\mathcal{S}\subset^{n}\times^{m}$, recovering a standard notion of merit function studied in the context of saddle-point optimization (Nemirovski et al., 2009). It is insightful to compare this bound side by side with the one we would get by running primal-dual stochastic gradient without regularization. By standard arguments (see, e.g., Zinkevich, 2003; Nemirovski et al., 2009), the following bound is easy to prove:
$\mathbb{E}\left[G(\bm{x}^{*};\bm{y}^{*})\right]\leq$ $\frac{\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]}{\eta_{x}T}+\frac{\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]}{\eta_{y}T}$
$+\frac{\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)\bm{y}_{t}+\widehat{\bm{b}}(t)\right\|_{2}^{2}\right]$
$+\frac{\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widehat{\bm{M}}(t)^{\mathsf{\scriptscriptstyle T}}\bm{x}_{t}-\widehat{\bm{c}}(t)\right\|_{2}^{2}\right].$
@@
At first, it may seem surprising that such an improvement is possible to achieve by such a simple regularization trick. To provide some insight about how regularization helps us achieve our goal, we provide the brief proof sketch of the above statement here (and defer the full proof to Appendix A.1).
@@
Consider $(\bm{x}^{*},\bm{y}^{*})\in^{m}\times^{n}$. As the first step, we introduce the notation $f^{\left(\text{reg}\right)}\left(\bm{x},\bm{y}\right)=f\left(\bm{x},\bm{y}\right)+\frac{\varrho_{x}}{2}\left\|\bm{x}-\bm{x}_{1}\right\|_{2}^{2}-\frac{\varrho_{y}}{2}\left\|\bm{y}-\bm{y}_{1}\right\|_{2}^{2}$ and rewrite the expected duality gap as
@@
These terms can then be controlled via the standard regret analysis of COMID due to Duchi et al. (2010), and an additional “ghost-iterate” trick to account for adaptive comparators (see Section 3 of Nemirovski et al., 2009). In particular, a few lines of calculations (along the lines of the online gradient descent analysis of Zinkevich, 2003) yield the following bound on the sum of the two regrets:
@@
$\qquad\leq\frac{\mathbb{E}\left[\left\|\bm{y}^{*}-\bm{y}_{1}\right\|_{2}^{2}\right]}{\eta_{y}T}+\frac{\eta_{y}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{y}(t)\right\|_{2}^{2}\right]$
$\qquad\quad+\frac{\mathbb{E}\left[\left\|\bm{x}^{*}-\bm{x}_{1}\right\|_{2}^{2}\right]}{\eta_{x}T}+\frac{\eta_{x}}{T}\sum_{t=1}^{T}\mathbb{E}\left[\left\|\widetilde{\bm{{g}}}_{x}(t)\right\|_{2}^{2}\right]$
@@
As can be seen from the proof sketch above, the role of the additional regularization term for the $x$-player is to eliminate the gradient norms appearing in the regret bound of the $y$-player. This effect kicks in once the regularization parameter $\varrho_{x}$ becomes large enough, so that the corresponding negative term in the regret bound of the first player can overpower the positive term appearing on the bound of the opposite player. The same story applies to the second player. Note that while the regularization pulls the iterates closer to the initial point $\bm{x}_{1},\bm{y}_{1}$, it does not explicitly guarantee that they remain uniformly bounded at all times $t$, and in fact such claim seems impossible to show in general due to the noise in the gradient estimates. Remarkably, the analysis above works seamlessly for noisy gradient estimates, even though the gradient noise can grow proportionally with the size of the iterates. Another technical challenge that the analysis needs to address is the potential data-dependence of the comparators $(\bm{x}^{*},\bm{y}^{*})$, which introduces a potential bias into the estimates of the gap function $G(\bm{x}^{*},\bm{y}^{*})$. We handle this bias by adapting an elegant “ghost-iterate” trick of Nemirovski et al. (2009), which we regard as an application of the technique of Rakhlin & Sridharan (2017) for controlling suprema of a collection of martingales via online learning.